OpenStudy (anonymous):

A formula is expressed as D=a(2+kt). Express k in terms of D, a and t.

OpenStudy (anonymous):

1) divide by \(a\) 2) subtract \(2\) 3) divide by \(k\)

OpenStudy (anonymous):

oops i meant 3) divide by \(t\)

OpenStudy (anonymous):

btw when i say "divide" i don't really mean "divide" i really mean "write it"

OpenStudy (anonymous):

\[ D=a(2+kt)\\ \frac{D}{a}=2+kt\] etc

OpenStudy (anonymous):

im working it right now :)

OpenStudy (anonymous):

i - the 2 and so i had D/a -2= kt

OpenStudy (anonymous):

when i divide the t how would i write that?

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

I do it a little different...but it should come out the same D = a(2 + kt) --- distribute through the parenthesis D = 2a + kta -- subtract 2a from both sides D - 2a = kta -- now divide by at (D - 2a) / (at) = k D/(at) - 2/t = k

OpenStudy (anonymous):

it says the answer is D-2a/at

OpenStudy (anonymous):

dont know how

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

the answer is in there

OpenStudy (midhun.madhu1987):